OpenStudy (anonymous):

Why is it easier to factor a trinomial when a = 1 than to factor a trinomial when a = something other than 1?

OpenStudy (anonymous):

Assuming that "a" refers to the variable in the following formulas for factoring trinomials: a³+ b³ = (a + b)(a² – ab + b²) a³ – b³ = (a – b)(a² + ab + b²) ___________________________________ When "a" =1, then not only will ALWAYS: a³ =1 ; AND: a² =1 ; BUT: As regards to "ab": "ab", (which equals "a*b") will away equal "b", (assuming b ≠ 0); since anything multiplied by "1" (except "zero"); equals that same number.

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

ya :)